OpenStudy (anonymous):

Does anyone here go to COnnections academy and can help me I will give medal and fan

OpenStudy (preetha):

Please ask your question. There are many who can help.

OpenStudy (anonymous):

I already did and No one would help me

OpenStudy (solomonzelman):

I go to Connections academy:) But, as Preetha says, ask your question. I'll try my best, and so will do many other people!

OpenStudy (anonymous):

Simplify the radical expression \[3\sqrt{5} + 6\sqrt{45}\]

OpenStudy (anonymous):

I have tried over and over again and I can't figure it out @SolomonZelman

OpenStudy (solomonzelman):

the problem is \(\LARGE\color{purple}{ \bf 3 \sqrt{5} +6\sqrt{ 45} }\) at first, simplify √45 \(\LARGE\color{purple}{ \bf \sqrt{ 45} = \sqrt{ 9 \times 5} = \sqrt{ 9} \times\sqrt{ 5}=3\sqrt{5}}\) Correct?

OpenStudy (anonymous):

That isn't one of the choices :/

OpenStudy (anonymous):

\[9\sqrt{50}\] \[21\sqrt{10}\] \[6\sqrt{96}\] \[21\sqrt{5}\]

OpenStudy (anonymous):

Those are the choices I got \[9\sqrt{50}\]

OpenStudy (solomonzelman):

I know this is not one of the choices. The thing I said before, I was trying to go through the process with you. What you got is incorrect, and you are clearly not following what I was trying to say over here.

OpenStudy (solomonzelman):

Do you see how I simplified √45 ?

OpenStudy (anonymous):

I understand that, But why do you have to change \[\sqrt{45}\]

OpenStudy (anonymous):

Yes but how does that equal\[3\sqrt{5}\]

OpenStudy (solomonzelman):

So that \(\large\color{black}{ \bf 3 \sqrt{5} }\) and √45 can be like terms, and so that I can add them.

OpenStudy (preetha):

Jessica, Solomon has shown you the first step. Now if you move sort(9) outside the sort, you'll get 3(sqrt5).

OpenStudy (solomonzelman):

\(\LARGE\color{purple}{ \bf \sqrt{ 45} = \sqrt{ 5 \times 9}}\) @JessicaBlakely CORRECT ?

OpenStudy (anonymous):

Yes I get it now, so now how do we add them?

OpenStudy (anonymous):

So now it looks like this correct? \[3\sqrt{5 + 6} \sqrt{9 * 5}\]

OpenStudy (solomonzelman):

Hold..

OpenStudy (solomonzelman):

\(\LARGE\color{purple}{ \bf \sqrt{ 45} = \sqrt{ 5 \times 9}}\) \(\LARGE\color{red}{ \bf \sqrt{ 5 \times 9}=\sqrt{ 9} \times \sqrt{ 5}}\) \(\LARGE\color{green}{ \bf \sqrt{ 9} \times \sqrt{ 5} = \sqrt{ 3^{2}} \times \sqrt{ 5} }\) \(\LARGE\color{royalblue}{ \bf \sqrt{ 3^{2}} \times \sqrt{ 5} =3\times \sqrt{ 5}=3\sqrt{ 5}}\) verify and make sure that you get each step.

OpenStudy (anonymous):

I do not get the green and blue part

OpenStudy (anonymous):

I do now! you put an exponent on the 3 to equal 9?

OpenStudy (solomonzelman):

yes!

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \bf \sqrt{ 45}=3\sqrt{ 5} }\) correct?

OpenStudy (anonymous):

Yes! so wouldnt that cancel out the other \[3\sqrt{5}\]

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \bf 6\sqrt{ 45}=6 \times(3\sqrt{ 5} )}\)

OpenStudy (solomonzelman):

right?

OpenStudy (anonymous):

Yes! and then we add 5?

OpenStudy (anonymous):

or we times 5 by \[3\sqrt{5}\]

OpenStudy (solomonzelman):

\(\LARGE \color{blue}{ \bf 3(6 \sqrt{ 5})=18\sqrt{ 5} }\) Now.... \(\LARGE \color{red}{ \bf 3 \sqrt{ 5}+6\sqrt{ 45} }\) is the same thing as saying.... \(\LARGE \color{blue}{ \bf 3 \sqrt{ 5}+18\sqrt{ 5} }\) \(\LARGE \color{blue}{ \bf 3 \sqrt{ 5}+18\sqrt{ 5}=(3+18)\sqrt{ 5}=21\sqrt{ 5} }\)